# INRetouch: Context Aware Implicit Neural Representation for Photography Retouching

### v2

@@ 6 More Visual Results
We show in Fig. C the qualitative results of various methods for the retouching transfer task. Our approach excels in accurately learning the edits from before-and-after image pairs, producing outputs that are not only more realistic but also better aligned with the intended edits. In contrast, other methods struggle to achieve similar fidelity, often resulting in noticeable artifacts and inconsistencies. This highlights the effectiveness and robustness of our method in capturing and applying complex retouching transformations.
Figure C: Comparison between different methods for our retouching transfer task
@@ 7 Presets for Our Dataset
To ensure the versatility and robustness of our dataset, we curated a diverse collection of varying presets, designed to simulate a wide range of editing styles and conditions. As shown in Fig. D, we apply some of these presets to a single natural image, showcasing the richness and variety inherent in the dataset. This comprehensive coverage not only highlights the adaptability of our approach to diverse editing scenarios but also establishes our dataset as a valuable resource for developing and evaluating methods capable of handling complex retouching tasks. Such diversity enables the models trained on our dataset to generalize effectively across different styles. We showcase the ability of our method to simulate these presets on videos in our accompanying video.
@@
Figure D: Visualization of the output of our different presets applied to a natural image (highlighted top left). Zoom in to see better.

### v3

@@ 4 More Visual Results
We show in Fig. D the qualitative results of various methods for the retouching transfer task. Our approach excels in accurately learning the edits from before-and-after image pairs, producing outputs that are not only more realistic but also better aligned with the intended edits. In contrast, other methods struggle to achieve similar fidelity, often resulting in noticeable artifacts and inconsistencies. This highlights the effectiveness and robustness of our method in capturing and applying complex retouching transformations.
Figure D: Comparison between different methods for our retouching transfer task
@@ 5 Presets for Our Dataset
To ensure the versatility and robustness of our dataset, we curated a diverse collection of varying presets, designed to simulate a wide range of editing styles and conditions. As shown in Fig. E, we apply some of these presets to a single natural image, showcasing the richness and variety inherent in the dataset. This comprehensive coverage not only highlights the adaptability of our approach to diverse editing scenarios but also establishes our dataset as a valuable resource for developing and evaluating methods capable of handling complex retouching tasks. Such diversity enables the models trained on our dataset to generalize effectively across different styles. We showcase the ability of our method to simulate these presets on videos in our accompanying video.
@@
Figure E: Visualization of the output of our different presets applied to a natural image (highlighted top left). Zoom in to see better.